• Find the area of the surface of the solid formed by the revolution the curve, x=a(θ - sinθ) and Y=a(1- cosθ) about x-axis (y=0).
1
Expert's answer
2021-06-09T11:15:02-0400

"A = 2\pi\int_{\alpha}^{\beta}x(θ) \sqrt{[x′(θ)]2+[y′)]2dθ}"

"Parameter\space θ \space varies\space from\space 0 \space to \space2π"

Derivatives:

"x\n\u2032\n(\n\u03b8\n)\n=\n(\n\u03b8\n\u2212\nsin\n\u03b8\n)\n\u2032\n=\n1\n\u2212\ncos\n\u03b8\n,"

"y\n\u2032\n(\n\u03b8\n)\n=\n(\n1\n\u2212\ncos\n\u03b8\n)\n\u2032\n=\nsin\n\u03b8\n,"

"[x′(θ)]^2+[y′(θ)]^2=(1−cosθ)^2+sin^2θ=1−2cosθ+cos^2 \theta+sin^2 \theta"

"=2-2cos \theta=4 sin ^2 \frac{\theta}{2}"

"A=2\pi\int_0^{2x} [(\theta-sin \theta).2sin \frac{\theta}{2}]d\theta=4 \pi[\int_0^{2x} \theta sin \frac{\theta }{2} d \theta- \int_0^{2x} sin \theta sin \frac{\theta}{2} d\theta]"

"I_2=\int_0^{2x} sin \theta sin \frac{\theta}{2}d \theta=\frac{4}{3} sin^3 \frac{\theta}{2} |_0^{2x}=0"

The area of the surface is ;

"=4x[I_1-I_2]"